# A Multi-Scale Feature Fusion Framework Integrating Frequency Domain and Cross-View Attention for Dual-View X-ray Security Inspections

### v2

@@ IV-A Dataset Description
TABLE I: Overview of Current Multi-View X-Ray Datasets
Dataset Positive Negative Categories Availability
DvXray [7] 5,000 11,000 15 Open
MV-Xray [24] 3,231 950 2 No
DBF [25] 2,528 0 4 No
This study utilizes the DvXray [7] dataset provided by Bowen Ma et al., aimed at advancing prohibited item detection in luggage security screening. DvXray is a large-scale dual-view X-ray image dataset comprising 32,000 images, including 5,000 pairs of positive samples and 11,000 pairs of negative samples. The dataset samples are sourced from real subway stations (approximately 66%) and manually designed luggage scanned using X-ray scanners (approximately 34%). It includes 15 categories of common prohibited items, such as firearms, knives, and explosives. All images are in PNG format with a resolution of 800$\times$600 pixels and are annotated at the image level.
As shown in Table I, DvXray is currently the only publicly available dual-view X-ray dataset. Other dual-view datasets are either not publicly released or do not match the scale and category coverage of DvXray. Therefore, DvXray was selected as the primary data source to ensure the reproducibility and reliability of the experiments.
@@ IV-B1 Experimental Environment
All models were initialized with ImageNet-1K pre-trained weights to leverage the knowledge learned from the large-scale dataset. Additionally, to enhance training efficiency and model stability, we applied Automatic Mixed Precision (AMP) [28] and Exponential Moving Average (EMA) techniques [29]. These strategies help to accelerate convergence and improve model accuracy by reducing memory usage and smoothing model updates during training.
@@ IV-B2 Loss function
$\text{L}(x,y)=-\frac{1}{C}\sum_{i=1}^{C}\left[y_{i}\cdot\log(\sigma(x_{i}))+(1-y_{i})\cdot\log(1-\sigma(x_{i}))\right]$ (31)
@@
IV-B3 mAP and AP Calculation
In multi-label multi-class classification tasks, AP (Average Precision) and mAP (mean Average Precision) are commonly used to evaluate model performance.
Average Precision (AP)
AP is calculated for each class by averaging the precision at different recall levels:
@@
where $P(k)$ is the precision at the $k$-th prediction, and $\Delta R(k)$ is the change in recall at that prediction.
Mean Average Precision (mAP)
mAP is the mean of AP values across all classes:
@@
where $C$ is the total number of classes. mAP gives an overall measure of model performance across all categories.
IV-C Comparison with Other Methods
TABLE II: Comparison of Model Performance Across Different Backbone
Model FLOPs Params Val_mAP Test_mAP
ResNet50-dual 10.801G 31.929M 0.8093 0.7895
+AHCR 12.405G 34.882M 0.8249 0.8073
+Ours 14.200G 48.302M 0.8533 0.8498
ResNeXt50_32x4d-dual 11.192G 23.041M 0.8696 0.8580
+AHCR 12.808G 34.354M 0.8721 0.8684
+Ours 14.604G 47.774M 0.8754 0.8761
RegNet_x_3.2gf-dual 8.414G 14.318M 0.8291 0.8240
+AHCR 8.717G 16.865M 0.8352 0.8286
+Ours 9.226G 20.089M 0.8389 0.8430
ConvNeXt_Tiny-dual 11.810G 30.194M 0.8823 0.8848
+AHCR 11.887G 29.438M 0.8971 0.8933
+Ours 12.435G 31.381M 0.9151 0.9098
To validate the adaptability and effectiveness of the proposed method across different backbone networks, we conducted comparative experiments using ResNet50 [30], ResNeXt50_32x4d [31], RegNet_x_3.2gf [32], and ConvNeXt_Tiny [33] as backbone models. Each backbone was tested with the baseline model (dual), a comparison method (AHCR), and the proposed approach (Ours). The results are presented in Table II.
Overall, the proposed method consistently outperforms both the baseline and AHCR in terms of Val_mAP and Test_mAP across all backbone networks, demonstrating its effectiveness in enhancing dual-view X-ray image classification and its strong generalization capability.
In addition, dual labeling based on the backbone model involves processing the OL and SD views through a shared-weight backbone, concatenating the resulting features, and subsequently passing them through $Conv_{1}$ before forwarding them to the classification head. Further details are available in our open-source code.
A detailed analysis is as follows:
1. Performance on ResNet50 Compared to the baseline model, the proposed method improves Val_mAP from 0.8093 to 0.8533 and Test_mAP from 0.7895 to 0.8498, significantly outperforming AHCR (0.8249 and 0.8073, respectively). This indicates that the proposed enhancements effectively exploit dual-view information, yielding substantial improvements even in a conventional CNN architecture.
2. Performance on ResNeXt50_32x4d Leveraging grouped convolutions, this architecture provides stronger feature extraction capabilities, leading to higher classification accuracy across all methods. The proposed approach achieves a Val_mAP of 0.8754, a 0.58% improvement over the baseline and a 0.33% improvement over AHCR, with similar gains observed in Test_mAP. This result suggests that the proposed method can further enhance performance even within more advanced feature extraction frameworks.
3. Performance on RegNet_x_3.2gf Designed for optimized parameter efficiency, this network maintains relatively low computational complexity while delivering competitive classification performance. The proposed method improves Val_mAP from 0.8291 to 0.8389 and Test_mAP from 0.8240 to 0.8430, demonstrating its robustness and adaptability.
4. Performance on ConvNeXt_Tiny Incorporating CNN and Transformer-inspired architectural principles, ConvNeXt provides superior feature representation. The proposed method achieves a Val_mAP of 0.9151, marking a 3.29% increase over the baseline and a 1.80% increase over AHCR, while also achieving a Test_mAP of 0.9098. These results confirm that the proposed enhancements generalize well across networks of varying depths and feature extraction paradigms.
IV-D Ablation Study Analysis
TABLE III: Ablation Study on Different Model Components
@@ IV-D Ablation Study Analysis
ResNet50-dual 10.801G 31.929M 0.8093
+CAFM 11.875G 42.979M 0.8252
+MSCFE 12.990G 37.227M 0.8336
+FDIM 12.957G 32.340M 0.8342
+MSCFE & FDIM 15.146G 37.637M 0.8304
+CAFM & MSCFE 14.064G 48.276M 0.8446
+CAFM & FDIM 14.031G 43.389M 0.8334
+ALL 16.220G 48.686M 0.8527
@@
To evaluate the impact of different functional modules on model performance, we conducted ablation experiments on the DvXray dataset. Previous research has demonstrated the superiority of dual-view imaging over single-view for mitigating object occlusion; therefore, ResNet50-dual was directly adopted as the baseline model. On this basis, various feature enhancement modules were progressively incorporated to investigate their contributions to classification performance. The experimental results are presented in Table III.
For individual module evaluation, CAFM improves the channel-wise feature representation, increasing Val_mAP from 0.8093 to 0.8252. While this confirms its effectiveness in feature fusion, the performance gain remains limited when used alone. MSCFE further enhances feature interactions between dual views, improving cross-view information utilization and achieving a Val_mAP of 0.8336, representing a 2.99% increase over the baseline. FDIM optimizes feature alignment across views by guiding feature fusion in the frequency domain, leading to a Val_mAP of 0.8342, demonstrating its effectiveness in reducing viewpoint discrepancies.
In module combination experiments, the integration of CAFM and MSCFE further improves Val_mAP to 0.8446, which is 1.1% higher than using MSCFE alone. This indicates that the synergy between channel attention and multi-scale interactive feature extraction enhances feature fusion. The combination of CAFM and FDIM achieves a Val_mAP of 0.8334, slightly lower than FDIM alone, suggesting that channel attention and frequency-domain alignment may introduce redundant information under certain conditions, affecting model convergence. The MSCFE and FDIM combination results in a Val_mAP of 0.8304, lower than their individual performances, which may be attributed to conflicts between multi-scale feature extraction and frequency-domain feature alignment, leading to feature distribution inconsistencies.
When all modules are integrated (+ALL), the model achieves the best performance with a Val_mAP of 0.8527, representing a 5.36% improvement over the baseline. This confirms that each module contributes to different aspects of feature representation and cross-view alignment, and their collaborative optimization further enhances classification performance. However, the results also indicate that the impact of module combinations varies, with MSCFE contributing the most among single modules, whereas FDIM exhibits instability when combined with other modules. Therefore, in practical applications, selecting appropriate module combinations based on specific scenarios and computational constraints is essential to achieving an optimal balance between efficiency and accuracy.
@@ V Conclusion
In this paper, we presented a novel multi-scale interactive feature fusion framework for dual-view X-ray security inspection. Our method enhances feature representation through frequency-domain interactions and multi-scale attention, significantly improving classification accuracy. Experimental results show that our approach consistently outperforms state-of-the-art methods across multiple backbone networks, demonstrating its robustness and adaptability. Future work will focus on further improving feature fusion techniques and exploring real-world applications for security screening.
Acknowledgment
This work was supported in part by a national science foundation, and in part by other supporting funds. The authors would like to express their gratitude to all anonymous reviewers for their valuable comments and suggestions, which have significantly improved the quality of this paper.

### v4

@@ IV-A Dataset Description
This study utilizes the DvXray[7] dataset provided by Bowen Ma et al., aimed at advancing prohibited item detection in luggage security screening. DvXray is a large-scale dual-view X-ray image dataset comprising 32,000 images, including 5,000 pairs of positive samples and 11,000 pairs of negative samples. The dataset samples are sourced from real subway stations (approximately 66%) and manually designed luggage scanned using X-ray baggage scanner (approximately 34%). It includes 15 categories of common prohibited items, such as firearms, knives, and explosives. All images are in PNG format with a resolution of 800$\times$600 pixels and are annotated at the image level.
DvXray is currently the only publicly available dual-view X-ray dataset. Other dual-view datasets are either not publicly released or do not match the scale and category coverage of DvXray. Therefore, DvXray was selected as the primary data source to ensure the reproducibility and reliability of the experiments.
@@
TABLE I: Comparison of Model Performance Across Different Backbones
Model FLOPs Params Val_mAP Test_mAP
ResNet50-Dual[32] 10.793G 23.539M 80.9% 80.5%
+AHCR[7] 12.409G 34.882M 83.7% 83.8%
+Ours 15.671G 35.611M 85.3% 85.3%
ResNeXt50_32x4d-Dual[33] 11.196G 23.011M 87.0% 86.3%
+AHCR[7] 12.812G 34.354M 87.2% 86.8%
+Ours 16.075G 35.082M 89.2% 89.2%
RegNet_x_3.2gf-Dual[34] 8.414G 14.303M 81.8% 81.0 %
+AHCR[7] 8.717G 16.865M 83.5% 82.9%
+Ours 9.240G 16.784M 86.9% 85.3%
ConvNeXt_Tiny-Dual[35] 11.659G 27.822M 88.2% 88.5%
+AHCR[7] 11.891G 29.438M 89.7% 89.3%
+Ours 12.356G 29.534M 91.5% 91.0%
@@ IV-B1 Experimental Environment
All models were initialized with ImageNet-1K pre-trained weights to leverage the knowledge learned from the large-scale dataset. Additionally, to enhance training efficiency and model stability, we applied Automatic Mixed Precision (AMP)[30] and Exponential Moving Average (EMA)[31]. These strategies help to accelerate convergence and improve model accuracy by reducing memory usage and smoothing model updates during training.
@@ IV-B2 Loss function
$\text{L}(x,y)=-\frac{\sum_{i=1}^{C}\left[y_{i}\cdot\log(\sigma(x_{i}))+(1-y_{i})\cdot\log(1-\sigma(x_{i}))\right]}{C}$ (17)
@@
IV-C Evaluation Metrics
In multi-label multi-class classification tasks, mAP (mean Average Precision) and AP (Average Precision) are commonly used evaluation metrics to measure the performance of classification models.
Average Precision (AP) Calculation
For each class, AP is calculated as follows:
@@ IV-C Evaluation Metrics
where $P(k)$ is the precision at the $k$-th prediction, $R(k)$ is the recall at the $k$-th prediction, and $\Delta R(k)$ is the change in recall at the $k$-th prediction.
The calculation steps are as follows: Precision: The ratio of true positive predictions to all predicted positives:
$P=\frac{TP}{TP+FP}$ (20)
where TP is true positive, and FP is false positive.
Recall: The ratio of true positive predictions to all actual positives:
$R=\frac{TP}{TP+FN}$ (21)
where FN is false negative.
AP is calculated by sorting predictions by confidence, computing precision and recall for each prediction, and then averaging over all predictions.
Mean Average Precision (mAP) Calculation
mAP is the average of AP across all classes. For $C$ classes, mAP is calculated as:
@@
where $\text{AP}_{c}$ is the average precision for class $c$, and $C$ is the total number of classes.
mAP reflects the overall performance of the model across multiple classes. In multi-label tasks, each image may have multiple labels, and mAP gives an overall performance measure by averaging the AP of all classes.
TABLE II: Ablation Study on Different Model Components
@@
ResNet50-Dual 10.793G 23.539M 80.9%
+CGFM 11.335G 32.480M 82.5%
+DVHEM 12.982G 26.194M 82.4%
+FDIM 12.941G 24.014M 81.4%
+DVHEM & FDIM 15.129G 26.671M 83.0%
+CGFM & DVHEM 13.524G 35.135M 84.5%
+CGFM & FDIM 13.483G 32.955M 83.3%
+ALL 15.671G 35.611M 85.3%
IV-D Comparison with Other Methods
To validate the adaptability and effectiveness of DAGNet across different backbone, we conducted comparative experiments using ResNet50[32], ResNeXt50_32x4d[33], RegNet_x_3.2gf[34], and ConvNeXt_Tiny[35] as backbone models. Each backbone was tested with the baseline model (DUAL), a comparison method (AHCR), and the proposed approach (Ours).
Overall, DAGNet consistently outperforms both the baseline and AHCR in terms of Val_mAP and Test_mAP across all backbone, demonstrating its effectiveness in enhancing dual-view X-ray image classification and its strong generalization capability.
Additionally, for the dual‐view baseline we follow the AHCR configuration: both the OL and SD)views are forwarded through a shared‐weight backbone network, and for each class the maximum activation across the two view‐specific outputs is selected as the final prediction. Further implementation details are available in our open‐source code.
As shown in Table I, DAGNet achieves 6.4%, 1.8%, 1.9%, and 2.5% improvements over the baseline models ResNet50, ResNeXt50_32x4d, RegNet_x_3.2gf, and ConvNeXt_Tiny in terms of test set mAP, respectively. and also demonstrates superior performance compared to AHCR. These results indicate that DAGNet can stably improve classification performance on different backbone, especially showing more obvious performance advantages on networks with more complex deep learning architectures, verifying the effectiveness and superiority of the method in the dual-view X-ray image classification task. In summary, DAGNet outperforms the baseline methods and AHCR on the selected common backbone, demonstrating the effectiveness and superiority of DAGNet in the dual-view X-ray image classification task.
IV-E Ablation Study Analysis
@@
To evaluate the impact of different functional modules on model performance, we conducted ablation experiments on the DvXray dataset. Previous research has demonstrated the superiority of dual-view imaging over single-view for mitigating object occlusion; therefore, ResNet50-dual was directly adopted as the baseline model. On this basis, various feature enhancement modules were progressively incorporated to investigate their contributions to classification performance. The experimental results are presented in Table II.
From the experimental results, it can be observed that the validation set mAP of the baseline model ResNet50-dual is 80.9%. When the CGFM module is added alone, the validation set mAP improves to 82.5%, indicating that the convolution-guided fusion module can significantly improve model performance by effectively highlighting key features; while the DVHEM module effectively utilizes the complementary information between dual perspectives, further improving the validation set mAP to 82.4%, an increase of 1.5% over the baseline model. The FDIM module also significantly improves the fusion effect of cross-view features through frequency domain feature interaction and alignment strategies, achieving a validation set mAP of 81.4%.
In terms of module combination, CGFM combined with DVHEM achieved a validation set mAP of 84.5%, indicating that the convolution-guided fusion module and cross-view multi-scale feature interaction can effectively collaborate to further improve the overall feature fusion effect. However, the combination of CGFM and FDIM performed slightly worse than using the FDIM module alone, suggesting that the convolutional-guided fusion module and frequency-domain alignment may be redundant under certain conditions, affecting the model’s effective convergence. However, when all modules are combined, the model achieves optimal performance, with a validation set mAP of 85.3%, representing an improvement of 4.4% over the baseline model. This highlights the distinct advantages of each module in feature representation and cross-view information alignment, and their synergistic optimization significantly enhances the model’s overall classification performance.
Furthermore, from the perspective of computational resources, although the model has increased in terms of parameter count and computational complexity, considering the significant performance improvement, this increase in computational overhead is reasonable and efficient. These results fully demonstrate the complementary and synergistic enhancement mechanisms of the three proposed modules in the feature extraction, feature interaction, and feature fusion stages, collectively constructing a complete and efficient dual-perspective feature processing framework.
@@ V Conclusion
In this paper, we presented a Dual-views attention-guided network model, DAGNet, for dual-view X-ray security inspection. By introducing three core modules—FDIM, DVHEM, and CGFM—it effectively addresses feature misalignment and information loss caused by view differences and object occlusion in dual-view X-ray images.
Each module has its own unique design: FDIM enhances the extraction of detailed features through frequency domain transformation and dynamic interaction of frequency domain information between perspectives; DVHEM achieves feature alignment between two perspectives and multi-scale feature enhancement through a cross-attention mechanism and a hierarchical attention guidance module; CGFM efficiently fuses dual-perspective features and suppresses redundant information through convolution attention guidance.
Experimental results demonstrate that the method significantly improves classification performance on multiple backbone, fully validating the effectiveness and necessity of the collaborative work of each module. Overall, this paper not only provides a new feature fusion approach for dual-view security inspection systems but also lays a solid foundation for future research on application and optimization in real-world security inspection scenarios.